OpenStudy (naveenbhatia1312):

Verify the identity. Step by Step Please. Need Help!!! cotx/1+cscx=cscx - 1/cotx

jhonyy9 (jhonyy9):

like a first step rewrite cot x = cos x/sin x and 1+csc x = ? or csc x -1 = ? try it in this way ?

OpenStudy (abmon98):

\[\cot(x)/1+\csc(x)=\csc(x)-1/\cot(x)\] \[\cot(x)(\csc(x)-1)/(\csc(x)-1)(\csc(x)+1)\] \[\sin ^2(x)+\cos^2(x)=1 \] \[1+\cot ^2(x)=\csc^2(x) \] \[\csc^2(x)-1=\cot^2(x)\] \[\cot (x)(\csc(x)-1)/\csc^2(x)-1\] \[\cot(x)(\csc(x)-1)/\cot ^2(x)\] \[\csc(x)-1/\cot(x)\]